Consider that in a population of Filipino adults ages 18 to 65, Body Mass Index (BMI) is normally distributed with a mean of 28 and a standard deviation of 6. What is the BMI mark of the bottom 20% of the distribution of this population? *


1
Expert's answer
2021-10-18T06:09:39-0400

Let "X=" BMI, "X\sim N(\mu, \sigma^2 )"

Given "\mu=28, \sigma=6"


"P(X< x)=P(Z<\dfrac{x-\mu}{\sigma})"

"=P(Z<\dfrac{x-28}{6})=0.2"

"\dfrac{x-28}{6}\approx-0.8416"

"x\approx22.95"